“It is important for democratic media advance that participants be alert to and take advantage of every technological innovation.” — E. S.Herman, \underline{Democratic Media}, Z Papers

ELECTRONIC ACTIVISM: Part I

By Harel Barzilai

PURPOSE:

This article is in response to the call for articles (“long-standing projects ... accomplishments, plans, problems, and reasons why new people ought to become involved,” Journal of the 5th Year, Feb. ’92 issue). It is also a response to Edward Herman’s article \underline{Democratic Media} in Z Papers, since

(1) Herman’s otherwise comprehensive article lacks substantial coverage of electronic activism and

electronic media; and:

(2) We would like to show (in the second of the article’s two parts) how electronic activism can address some of the key problems and issues Herman raised as essential to “democratizing the media”, and how it can help overcome, more generally, obstacles to effective activism.

The article covers:

(A) WHAT’S OUT THERE: A (Partial) Overview of the Electronic Networks; (B) OUR PROJECT: Describing the projects we are working on; (C) HOW Electronic Activism can address problems raised by Herman, and how Electronic Activism

can enhance activism in general; (D) THINGS we need help with now, why people should get online and become electronic activists,

and possible future projects.

(A) and (B) are integrated into a “guided-tour of the electronic world” in the first part of the article, while (C) and (D) are dealt with in the second part of the article, which also sheds more light on (B).

OVERVIEW

The first part of this article is essentially a “crash course” on the electronic networks — more accurately, on some of the main ones, since there is a whole “electronic world” out there the description of which would take entire books. We will look at several facets of that world, and then, alongside each one, offer an example showing how it applies to activism. We will be going from simpler to more elaborate forms of electronic activism; in so doing, we will be tracing the evolution of the projects described more fully in the second part, which grew out of a series of earlier and simpler experiments in electronic activism.

The second part of the article describes the specific project a number of us are working on, including what we’ve accomplished, the future potential, and some problem areas and places where we need help.

⋆ ⋆ ⋆  ⋆ ⋆ ⋆  ⋆ ⋆ ⋆

IT WAS BACK IN ’86, when I was a first-year student at Dartmouth, that I started “playing” online, sending and receiving electronic mail (“email”, pronounced “E-mail”) — mainly between Dartmouth and Harvard University where a similarly inclined friend was enrolled. Although email was not new then, it was very exciting to be able to correspond with a friend miles away in this manner — without using stamps and envelopes, or spending money on long-distance phone calls.

Back then, the “transmission times,” or how long it took from the moment you told the computer to send your message to when your recipient’s machine received it, usually took several hours, and sometimes one and a half or even two days. Still, this was better than what the “online” community oftens refers to as “US mail” or “paper-mail” — a common derisive pun being to replace the “m” with an “n”: “I will send that by USnail” — that incredibly slow medium taking two or three days to arrive at its destination, or “Snail-mail.”

Today, email is an essential tool in a world-wide community of people and information-exchange. Every day, this community uses email to exchange countless pieces of information — not just letters, but articles, newsletters, and so on; other components in this world of telecommunications are “posts” on electronic bulletin board systems (BBSs), and even digitally encoded graphics and video images.

Email today is also generally much faster; at least, on the primary electronic networks (for example, the Internet) which are accessible from most major universities and many businesses. In fact, sending and receiving email is often “instantaneous,” taking under a second, and among the “wellconnected” networks, usually within a minute or so, even between locals thousands of miles apart.

As a test, I just logged into my California-based Peacenet account, and sent an electronic “test message” to my Cornell account where I am composing this article. It then took me seven seconds to return to my Cornell account and check my mail, by which time the message was there. This is what is meant by “essentially instantaneous” — as instantaneous as one is able to notice.^{1}

Many of us on the “Nets” — whether electronic activists, hackers, hobbyists, or just electronic pen-pals — have corresponded with people in Europe, Australia and New Zealand, and Japan, as well as Canada and Latin America^{2}, and transmission times between such remote locations can be very short as well, when messages are not sent during times of “peak load.” The “Culture Shock,” in the pleasant sense of the word, I can only compare with the wonderful experience, following my fairly insular high-school years, of going to college and meeting so many people from all over the world — including anti-apartheid activists from South Africa, friends I made from India, and others.

Email is — or rather, can be — a very intimate medium. I will not try to convey how or why this is so, leaving the discovery to be made by those who venture online. However, it is worth noting here that this is so, perhaps, because of and not in spite of the anonymity the medium affords — your race, sexual orientation, whether you are tall or short, how much you weigh, the way you dress, how often you shower, etc (and, if you have a name like mine, your gender) are not very obvious to the people with whom you communicate, nor are these attributes of others obvious to you. Ideas, humor, personality — whatever the participants choose to reveal of themselves, and not more — is what comes through.^{3}

1Peacenet and Econet are run by the Institute for Global Communication, and are described in more detail below. 2Africa and other parts of the third world are also getting online, including NGOs (Non-Governmental Organizations), in particular, in preparation for the UNCED (United Nations Conference on Environment and Development) conference this summer in Brazil. Microcomputed-based “FIDO” systems like NGONET and ESANET (Eastern and Southern African Network) are linked to the nodes of the Association for Progressive Communication (APC). APC is described below following “IGC — Networks for Activists”

See “Low cost global electronic communications networks for Africa” by Mike Jensen and Geoff Sears (Director of IGC), posted, Jan 19, 1992, to Peacenet’s “reg.africa” conference. 3Various visual aides have been developed to convey *italics* and _underlined_or >>EMPHASIZED<< words you can even <--- POINT at. On the Usenet (see below), there are the famous “smileys” conveying humor, anger, surprise, etc, and based on modifications of the sideways smiling-face comprised of three characters: (-:

How E-Mail Works:

I don’t know exactly how email works — any more than most people who don’t have a degree in electronic engineering know the precise workings of their television set — we will examine the analog of “How to use your Television set”

It should first be noted that there are many different mail-sending and mail-reading utility programs available, depending on the system you use. But generally, sending email is as simple as putting an address on an envelope. If you know a valid email address for your intended recipient (there can be more than one), you just type it into a line starting with a phrase like To: in your mail-sending program. The text of the letter is then either typed in or else “pasted” or “included” in from a pre-written file, and then the mailer program is given a command telling it: “send away!”

These email addresses are not very different from ordinary addresses, which start with an identification of the individual (their name), followed by information about the person’s location, starting with the narrow (street address) and proceeding to broader and broader “locators” (town/city, state, and zip). The same is done with email, with a few minor differences.

The first difference is that the entire address is put in one line; in fact, one expression. In this expression, the person’s account-name (their “user ID” on the computer they use) takes the place of the recipient’s name which would be written on the first line on the envelope, followed by an at-sign (@), and then the rest of the address, with dots (.) separating the parts comprising the “domain” or main part of the address.

For example, the email address of an online friend and fellow electronic activist, Rich Winkel at Missouri, is:

                       rich@pencil.cs.missouri.edu

Here rich is the account name Rich has on a computer nicknamed pencil at the Computer Science (cs) department at the University of Missouri, with edu (EDU-cational) being standard for InterNet addresses of colleges and universities.^{4}

Sometimes “smart” machines simplify addressing; for example, you don’t need to know the names of the machines — Sun workstations in this case — in the math department here at Cornell to send me email; rather than using:

               harelb@mssun7.msi.cornell.edu
      or:  harelb@wright.msi.cornell.edu

(“mssun7” and “wright” being among the “names” of these Suns)

You can email me at:

               harelb@math.cornell.edu

which a machine translates into the full address.

Some institutions have gone further. Dartmouth, for example, has a complete online “Dartmouth Name Directory” (DND) of all their students, faculty, and staff. The Dartmouth computer facilities receiving email use the DND to figure out how to reach the person you are trying to send email to.

The format used is “name@dartmouth.edu” — for example:

               john.w.lamperti@dartmouth.edu

In fact, less information can be supplied so long as it uniquely identifies one individual — thus lamperti@dartmouth.edu worked when this fellow electronic activist (and former math professor of mine) was the only “Lamperti” at Dartmouth; when professor Lamperti’s son was involved with Dartmouth (teaching rock-climbing), John could be reached by supplying the Dartmouth computers with an extra “j” to uniquely identify him: j.lamperti@dartmouth.edu

Not unlike a person would, the computer figured out from the information given which unique person “j lamperti” on the DND I had in mind.

This DND system is also set up so that if the recipient does not have an email account, the DND prints out the received email letter and delivers it to that person’s student or departmental mailbox.

4InterNet is sometimes confused with the UseNet, which is described below; the Internet is a major global network, parts of which are subsidized by various governments. It carries many kinds of traffic, including but not limited to Usenet BBSs. Usenet is also carried by other networks, for example, Bitnet.

Using email and the Usenet for Electronic Activism:

The origins of our current project start with the murder of the Jesuit priests in El Salvador in November of 1989, with my ordering Amnesty’s 1988 report, \underline{El Salvador: “Death Squads” — A Govern-} \underline{ment Strategy}, and with the outrage I felt upon discovering how clear and explicit the documentation had been, by none other than A.I., that the so-called “Death-Squads” — these shady right-wing extremist groups as the nooz-media invariably portrayed them — were in fact army/government run. Something had to be done to educate people about this, to use the tragic occasion to spread the word and put an end to Washington’s complicity (often much more than that) once and for all.

The email contacts I had were primarily people I had met by using the Usenet network of Electronic Bulletin Boards, or “newsgroups” as they are called (the Usenet is described in more detail below). By using newsgroups on topics such as “Activism”, “Latin-America,” and “Human Rights,” information could be exchanged, like-minded people met, and “online discussion” as well as planning could take place in the newsgroups and by email.

It may not be clear, a priori, that “electronic” planning and coordination of real-life projects (activist or otherwise) is necessarily an improvement over the exclusive use of the traditional methods (telephone, face-to-face mettings), or even that it is workable. However, so it proved in the present example, and it is my hope that the “time-line” of the projects we are tracing will provide concrete settings for the abstract — but very real — recurring themes:

• the money and especially the time saved;

• the ability to coordinate the activities of more people at more remote locations more easily;

• and the tools allowing for information-access and information-processing power on a scale which is a quantum leap ahead of the strictly non-electronic, traditional methods.

The project we organized was to send copies of the A.I. report to all U.S. Senators along with a signed letter demanding a change in U.S. policy, with copies going to the major media, stating also that this key Amnesty report had been made available to Congress, and what the report said. Using the electronic networks, and with two primary co-organizers at Dartmouth College (John Lamperti) and the University of Chicago (Mary Pugh), we were able to raise almost all the funds needed to buy and send the reports through online contacts.

Electronic mail was used to joinly revise the Statement to be sent to the Senators and the media, by rapidly swapping back and forth by email, half a dozen times, the latest draft. And, by “advertising” on the Usenet’s newsgroups, we were able to get the Statement signed by dozens of people from over twenty states. We were then able to have the statement signed by Archbishop Gumbleton of Detroit, one of the few co-signers^{5} whom we contacted by phone rather than email.

An overview of the Usenet

The Usenet is an interactive bulletin board system currently read by more than 2 million people worldwide, with some 10 million email accounts having access or potential access. (Electronic surveying devices keep these and other statistics which are cited below.)

The Usenet “system” is often referred to, with some merit, as an “anarchy” — in the political, not derisive/chaotic sense, although given the vast amount of people, data, and interactions in question (see below), “chaos” is unavoidable, and it is merely a question of how “well-managed” the chaos is.^{6}

5The co-signers gave their consent to add their name to the Statement, rather than signing one physically circulated copy of the Statement. 6To put this in perspective, it should be noted that there is substantially more “chaos”, or “random information” — really, huge, convoluted webs of interactions and cause/effect — in the non-electronic (“real”) world; it is merely that less of it is kept track of (e.g. how many items of “junk mail” do you get on average during the month of March? What are all the phone conversations you’ve had this month in which you mentioned the word “green”?) in ways the average person can access, so we ignore them and they join the informational “background noise” in our minds.

As one user put it, the Usenet is a “logical network” not a physical network (such as Internet)^{7}. In essence, informational “traffic” is passed along between the main Usenet “nodes” and thus the articles (as items posted to the newsgroups are called) get “propagated” across the system. In the U.S., anyone with a capable machine and the proper software can get a Usenet “feed” and thus become part of the Usenet.

Few of us can afford this means of access to the Usenet^{8}, however, and two of the most common ways to gain access is by means of a computer account at a major university or through where one works, by which students, faculty, and staff, and employees with an account can get on the net and read the Usenet newsgroups. From computers running the Unix operating system, news-readers such as “readnews,” “rn” and “nn” are available^{9}

From a Vax system, another popular system at universities, VMSNEWS is available. There is usually online help available — from unix, for example, type man rn to get the MANual description for that newsreader program; or e.g. help vmsnews on Vaxs — or else one can ask the local system operator for help.

If you cannot gain access to the Usenet from a university or your company, and if you have a computer and modem at home, there are public-access unix systems, or “Nixpubs,” which you can dial into using a modem. A monthly list of Nixpubs is posted to Usenet’s alt.bbs newsgroup, and many of these machines, some of which are freely accessible, carry Usenet “news” (the flow of “newsgroup” postings). In fact, by the time this article is published, I hope to have my mother in Massachusetts online thanks to these nixpubs, and thanks to tips about where to find good cheap modems which the readers of alt.bbs were kind enough to provide.

⋆ ⋆ ⋆  ⋆ ⋆ ⋆  ⋆ ⋆ ⋆

The individual sites which carry Usenet news are completely independent:

Every administrator controls his own site. No one has any real control over any site but his own.

The administrator gets her power from the owner of the system she administers. As long as her job performance pleases the owner, she can do whatever she pleases, up to and including cutting off Usenet entirely.^{10}

according to the official documents, insofar as anything on Usenet can be “official;” even in the one domain where Usenet functions as a “democracy” rather than “anarchy,” the voting procedure for the creation of a newsgroup (see below), one merely has The Official Guidelines.

The Usenet is also not “software” — there are dozens of software packages in use today for reading and transmitting Usenet articles. It is for these and other reasons that in describing Usenet as an electronic bulletin board “system” that word belongs in quotes, and why the official “What Is Usenet” description in the newsgroup news.newusers.questions spends most of its time stating what the Usenet is not, and finally contents itself with the seemingly all-purpose description:

Usenet is a set of people who exchange articles tagged with one or more universally-recognized labels, called “newsgroups.”

7Zbigniew J. Pasek, University of Michigan, Ann Arbor, compilation of Electronic Resources, Feb. ’92. 8The costs are not entirely prohibitive: the Usenet article describing What Is the Usenet (in the official monthly post to Usenet’s news.newusers.questions newsgroup) states that “Computers capable of taking Usenet feeds are down in the $500 range now, and UNIX-capable boxes are going for under $2000, and there are at least two UNIX lookalikes in the $100 price range.” 9“nn” standing for “No News” is Good News — see below regarding the quantity of “news traffic” on the Usenet. 10CAVEAT: You should not interpret this gender-neutral language in the official administrative posts to mean there isn’t any sexism on the Usenet — not to mention racism, homophobia, and some holocaust revisionists to boot.

These people reside at universities, research institutions, government agencies (e.g., one sees addresses like joe-shmoe@nasa.gov) as well as companies big and small (email addresses often ending in .com), and even high schools and private homes.

The Usenet Newsgroups

A typical Usenet newsgroup name, rec.sports.baseball, reveals two things: words/identifiers are separated by periods (pronounced “dot,” as in “rec[-reation] dot sports dot base-ball”); and that the name starts with a broad, general category, then a more specific identifier, and continuing in this manner if necessary.

The Usenet is organized in a logical hierarchy, by means of seven broad categories, and the first word in a “mainstream” Usenet newsgroup name is one of the following:

• news for things having to do with the Usenet itself; for example: news.newusers.questions or

      news.lists;

• comp, for example comp.sys.amiga, is all about computers;

    • rec is for “recreation” as in rec.sports.baseball, rec.games.chess, rec.music.classical,
      rec.food.veg, and so on;

• soc is for things “social” as in soc.women, soc.singles, soc.motss (“motss” is a Usenet-ism for Member Of The Same Sex; also in use are MOTOS (“opposite”) and MOTAS (“appropriate”)). there is also a whole culture sub-hierarchy — “soc.culture.*” where “*” can be celtic, or

      african, jewish, arab, japan, turkish, etc, etc;

• talk includes “high-noise” groups for “discussion” like talk.politics.misc, talk.abortion, and talk.guns (Go NRA!); few converts are made in these newsgroups’ debates, and many “flames” are exchanged between hostile parties, to cite another Usenet term;

• sci is for scientific or science-related discussions, e.g. sci.space, sci.math, sci.med.aids;

• misc includes “everything else” from misc.kids and misc.taxes to our own activist newsgroup, misc.activism.progressive (see below), which is specifically designed to be an informational resource, not a discussion (“talk”) group (for which we have an electronic mailing list).

There is also a hierarchy for unofficial alt-ernative newsgroups like alt.co-ops, alt.atheism, or alt.angst — no vote needs to be taken for the creation of a Usenet “alt” group, although the propagation of such newsgroups to other machines is more limited — administrators sometimes cut off all alt groups to avoid newsgroups like alt.sex and the rest of the alt.sex.* hierarchy. This happens despite the fact that alt.sex is usually at or near the top in the Usenet readership ratings — in February, for example, it had roughly 260,000 readers; that month, misc.jobs.offered was second (surprise, surprise) with 210,000 readers, and rec.humor.funny third with 180,000 readers — this according to the Usenet’s automated electronic “Nielsen”-type ratings system, the results being posted monthly to news.lists It should be noted that alt.sex is fairly tame by Usenet standards; in alt.sex.bondage, for example (still not the most daring newsgroup, and, interestingly, on a higher level of maturity than alt.sex) you can tune in to (or join) women and men discussing different types of knots, and the safe way to use gags, practice bondage, etc, with their “SO’s” (Significant Others).

Explosive Growth:

According to the newsgroup news.lists, in a two-week period in February, 195,672 articles, totaling some 388 Mega-bytes^{11}, were submitted from over eighteen thousand sites, by some 50,000

11A Mega-byte (Mbyte) is 1,024 Kilo-bytes, or “K” — for comparison, parts I and II of Noam Chomsky’s \underline{The Victors} which we have online, are about 70 K each. Thus the equivalent of roughly 400 such articles are posted each day — the content and overall “info to noise ratio,” however, are another matter.

If the “headers” of the articles are included, the figure comes to 476 Mbytes over the two weeks. users (or only about one in forty news readers) and to over 2,000 different newsgroups, totaling an average of 27.7 Mega-bytes (Mbytes) per day^{12}

The explosive growth of the Usenet is such that percentage growth is measured monthly — for example, below are the data from two readings at the end of last year (Note the growth approaching 40% in Mbytes-per-day between October of ’91 and the quoted figure for Feb. ’92):

                          91/10/23        91/11/27      growth
    Articles               142161          163345         15%
    Mbytes/day             20.124          22.330         11%
    Sites posting          14572           15860           9%
    Users posting          38388           43955          15%
    Newsgroups             1758            1891            8%

In fact, based on more detailed statistical analyses drawn from the data automatically collected by different Usenet machines, the estimate for August of 1995 volume is 365 Mbytes per day; for August of 2000: 12,154 Mbytes per day if the growth patterns continue.^{13}

In translation, on any given day of the week, many books’ worth of material is posted (many dozen books in 1995 if that rate is realized) on the Usenet alone, which does not count other networks, or the several thousand electronic mailing lists in operation.

Clearly, tools for narrowing down from this flood of information to what you are looking for, and making sense of that information, are crucial. Fortunately, the same technologies which opened this Pandora’s Box and its floods of information, can be used to design and implement exactly such tools, and the tools available today (“newsreaders” editors, etc) make the task far less formidable that this introduction, upon first reading, might suggest.

Electronic Mailing Lists and “Listservs”

Following the El Salvador project, I organized an (electronic) mailing-list of the participants, over which information and updates could be exchanged and through which long-term discussion and support could take place. This list, which was known as the Activists Mailing List (AML), rapidly grew as the word was spread by email and over the Usenet. Fortunately, by the time there list approached 100 members and was becoming difficult to manage “by hand,” one of AML members and original participants in the El Salvador project, Rich Winkel in Missouri, offered to put the AML on a “listserver”

12

                "Total traffic through uunet for the last 2 weeks"
                Message-ID: <1992Feb12.165056.13940@uunet.uu.net>
                Date: 12 Feb 92 16:50:56 GMT
                Sender: rick@uunet.uu.net (News Statistics)
    195672 articles, totaling 387.926123 Mbytes (476.163834 including headers),
    were submitted from 18092 different Usenet sites by 49442 different
    users to 2045 different newsgroups for an average of 27.709009 Mbytes
    (34.011702 including headers) per day.

13

    Brian Smithson (brian@csd.mot.com) in:
    Re-post of Usenet stats (LONG; was: Re USENET ANALYSIS REPORTs...)
    Message-ID: <4812@motcsd.csd.mot.com>

(courtesy of his system administrator) and thus came into being the ACTIV-L “Listserv”-based electronic mailing list.

The LISTSERV software manages an electronic mailing-list by creating an email address like ACTIV-L@UMCVMB.BITNET such that all email sent to that address is forwarded to everyone else who is subscribed to the mailing list. Listserv maintains and updates a list of all current subscribers, and “understands” simple commands such as JOIN ACTIV-L (which are emailed to it at LISTSERV@UMCVMB.BITNET) and by which people can have their address added to or removed from the mailing list (were it so easy to have one’s name removed from the “paper-mail” mailing lists!)

By maintaining a list of everyone currently subscribed, the Listserv eliminates the need for each person on the list to keep and update their own private list of current list members and to list for their computer’s mailer-program all the corresponding email addresses each time s/he wishes to post to the list.

There are today literally thousands of listserv-type electronic mailing lists, from AMALGAM (“Dental

AMALGAM and MERCURY Poisoning”) to YACHT-L (“YACHT-L@GREARN The Yachting Sailing and ama-

teur BoatBuilding”) and almost everything in between, for example:

         MAPS-L   MAPS-L@UGA.BITNET        Maps and Air Photo Systems Forum
         MEDIA-L  MEDIA-L@BINGVMB.BITNET   Media in Education
         MEDIMAGE MEDIMAGE@POLYGRAF.BITNET Medical Imaging Discussion List

Which brings us to another example of things Listserv is useful for. Since there are so many such lists, and and since the list of such lists changes increasingly rapidly as new lists are created, there is a mailing list, serviced by a Listserv, which sends out to it’s subscribers an announcement each time a new list is created summarizing what it is about, so you can decide if you want to join it, for example, something like:

        New list: PORTLVEG-L, issues of concern to Portuguese left-handed
        vegetarians. To Sign up, send email message ``SUB PORTLVEG-L'' to
        LISTSERV@STRANGE-LISTS.BITNET

This listserver-for-listservers does something else, too: it maintains an up to date list of all these mailing lists, adding new entries as lists are created, and removing names of archaic lists which “die out” as a result of lack of interest. You can give it a command to send you at any given time the most up to date list of mailing lists. Warning: the last time I sent out for this list I received a list of over 2800 mailing lists. However, a good unix editor like “emacs” (pronounced “E-macks”) has no trouble editing the such a big file, nor, for that matter, and almost instantly, finding in it all lines which have “dog” or “eco” or “fish” in them, allowing you to try out the handful of lists you might be interested in.

Many of the mailing lists (and Usenet newsgroups) are concerned with topics which may seem frivolous next to activism, but there are many useful non-activist forums where information can be shared or help solicited (taxes, home-repair, etc), as well as lists providing intimate support forums which the lists members might not have easy access to elsewhere (former abused-children, the disabled, etc).

⋆ ⋆ ⋆  ⋆ ⋆ ⋆  ⋆ ⋆ ⋆

The successes as well as frustrations and shortcomings of the El Salvador project led me to write up a proposal for a big project, called PACH (pronounced “patch”, and standing for Progressive Alliance Clearinghouse), which is the primary subject of the second half of this article.

At about the same time the Activists Mailing List “became” ACTIV-L, an ACTIV-L subscriber who also had an account with Peacenet (described in the following section), Andy Lang, forwarded my PACH files which I posted to ACTIV-L to then-Peacenet Director Howard Frederick.

Dr. Frederick was impressed enough to write back to offer a complimentary Peacenet account from which information and articles from the Peacenet and Econet systems could be forwarded on a regular basis to ACTIV-L (Rich Winkel, who set up the Listserver at Missouri for ACTIV-L, had been forwarding items occasionally from Peacenet for some time).

The Peacenet and Econet systems are part of broader nation-wide and world-wide networks, which are described below.

IGC — Networks for Activists

Peacenet and Econet are sister-networks dedicated solely to promoting peace, human-rights, social justice, and environmental issues through education and by providing resources for activists.

Jointly run by the Institute for Global Communications (IGC), a division of the non-profit Tides Foundation, based in San Francisco, these networks allow users to gather information from and post items to over 900 IGC newsgroups (called “conferences”) on topics ranging from Central and Latin America, Disarmament, Economics, Energy, and the Environment ... to Labor, the Media, the Middle East ... on through Toxics and Waste, Women, and Youth. There are also special conferences listing calendars of events, announcements, and general references and resources (not unlike “Zaps” online), and besides these conferences dedicated to news and information, there are others which serve as forums for discussions and for the planning of activism and long-term coordination; for example, Amnesty International has a whole array of “private” (internal) Peacenet conferences which it uses to coordinate its work).

IGC also offers its subscribers an email-to-fax link by which faxes can be sent for as little as 25 cents per page by addressing to the “email” address fax:1CCCNNNNNNN where 1-(CCC)-NNN-NNNN is the fax machine’s number. A similar service is offered for email-to-telex-machines, as well as telexmachine-to-email (there is no fax-to-email service as yet, although improved scanner technology and a better funded IGC may bring that about some day). IGC “News” conferences include wire services that are online, as well as legislative bulletins, newsletters, and even magazines that are online; for example, Propaganda Review is now entirely online.

IGC has over 7,000 users in the U.S., but deserves to be far better known, offering resources on a scale that would have been unimaginable before the advent of electronic networks.

One thing lacking in the unix-based IGC accounts is the power that comes with an individual unix account. However, if you have access to such a machine, through a university for example, IGC offers a “telnet” login as an alternative to modem dial-ins (“telnet” is a unix program allowing log-ins to remote machines) Telnetting into IGC, which is more convenient for the user, is actually preferred by IGC as it adds less to the load on their systems, and is offered on a discounted “off-peak” basis 24 hours a day, at $3/hour.

For me, this has meant “the best of both worlds,” using the emacs editor’s built-in telnet feature which provides a nice interface with IGC. Namely, emacs is designed for multi-tasking, whereby you edit several files at the same time, each residing in one of several buffers among which you can instantly switch. From emacs you can also send mail, receive mail, manipulate entire directories very nicely with its directory-editor dired and so on, each of these functions in a different buffer.

In this manner, IGC conferences can be read in one buffer while reading Internet mail, sending mail, etc, in the other buffers. An interesting Peacenet or Econet article can then be saved in my private directory, emailed, saved on a remote machine, or posted to the Usenet, each with a few keystrokes, by using one of the customizable emacs “macros”^{14}

To briefly survey the vast array of information on the IGC networks:

14Emacs’ customizability and expandability is legendary. I will say a little more in part II about the radical software-for-the-people project, the GNU project, of the Free Software Foundation — GNU Emacs being one result of this project. For example, emacs’ telnet facility was designed for an interface different from IGC’s, making the cursor “jump up” annoyingly; within a day or two of my posting to Usenet’s gnu.emacs.help about this, I was emailed a “patch” to modify (customize) emacs’s telnetting, nicely fixing the emacs-IGC telnet link.

Online news wires at IGC include Inter Press Service (IPS), ranked by UNESCO as the world’s fifth largest news agency. IPS is a non-profit cooperative and the leading Third World news agency, and provides several dozen stories per day through conferences (which typically hold roughly the 800 most recent stories at any given time). Other examples include the Environment News Service, the world’s only news service devoted exclusively to the environment; the U.N. Information Centre Newswire; the Harare-based Africa Information Afrique; and the Women’s Feature Service, based in New Delhi and New York.

U.S.-based news services include the Alternet News Service, a news and feature syndicate for the alternative press which also uses Peacenet as a medium for electronic distribution. Based in Washington, D.C., it was formed by alternative journalists to circumvent the corporate monopoly on information, and currently has one hundred newspaper clients in the U.S., distributing 20–25 stories per week. Another example is the San Francisco-based Pacific News Service, one of the world’s top alternative news agencies.

BOX: One of PeaceNet’s “INFO-COMPASS” login banners

      Welcome to PeaceNet / Bienvenido a PeaceNet.  For help, type "?"
                 <<<INFO-COMPASS for March 20-22, 1992>>>
Matti Peled Tour [Israeli Army Officer+Peace Activist]       mideast.action
                         ++++++++++++++++++++++++
COUNTERSPIN: FAIR Campaign Memo                                media.issues
Earth Summit: Talks Fatally Flawed+NGOs Denounce U.S.        ips.englibrary
East Timor Peace Ship/Rainbow Warrior+French Testing      nonviolent.action
Listing of Imprisoned Gulf Resisters                         military.draft
Madison Resolution on Cuba                                         reg.cuba
New World Order Dictionary!                                     pn.opinions
Radio station bans the name "Redskins"                        gen.nativenet
STOCKWELL: CIA's Secret Wars                      misc.activism.progressive
                    -==>> New Publications Online <<==-
RESULTS               hunger.general    PEOPLE'S TRIBUNE     gen.newsletter
FRONTLINE          nonviolent.action    MEXICO UPDATE        carnet.mexnews
PEACE NEWS                  wri.news    IPB NEWS                   ipb.news
MEDNEWS    misc.activism.progressive    INSIGHT FEATURES      nfd.ifeatures

You have new conf entries in pn.alerts pn.announcements pn.calendar ...
PeaceNet Commands:  (c)onferences (h)elp (m)ail (d)ata (u)sers (s)etup (bye)
?
-----------------------------------------------------------------------------

Also online are press releases by SANE/Freeze, Greenpeace, AFSC, and others; newsletters and bulletins, for example the Environmental and Energy Study Institute’s (EESI) “Weekly Bulletin”; and information about bills pending on Capitol Hill — for example, fcnl.updates contains the Friends Committee on National Legislation’s weekly legislative alert.

I cannot come close to covering the range and diversity of online resources on IGC, but to cite one more example, the “mosconews” conferences contain articles from Moscow News, an independent weekly newspaper founded in 1930 and distributed in more than 140 countries, and which is independent from all state and party structures in the former USSR.

IGC subscribers also have access to the several thousand Usenet newsgroups.

⋆ ⋆ ⋆  ⋆ ⋆ ⋆  ⋆ ⋆ ⋆

A.P.C.

IGC is a founding member of the Association for Progressive Communication (APC), a worldwide umbrella group of progressive networks dedicated to peace, human rights, and the environment. The APC Member Networks are: AlterNex (Brazil); Chasque (Uruguay); ComLink (Germany — note the pun); GlasNet (Russia); GreenNet (England); IGC (USA); Nicarao (Nicaragua); NordNet (Sweden); Pegasus (Australia); and Web (Canada). Membership is pending for EcuaNex (Ecuador) and Lega per L’Ambiente (Italy).

All APC partners are independent organizations, and retain full control over their network, paying fees to a fund to further the global spread of the network.

While there are 10 countries with member-networks, people in over 90 different countries call those machines; for example, many people in Africa and Europe call GreenNet. In all, there are some 15,000 APC users.

The subscribers of these networks can exchange electronic mail not only with other subscribers, but also with people on other networks, including Internet, Bitnet, CompuServe, the WELL, and MCI mail. The APC networks also share a more intimate link among themselves, by directly exchanging daily the information posted to their newsgroups, with a 2-way information “feed” between the different APC machines.

For more information about Peacenet or Econet, write to: IGC, 18 De Boom Street, San Francisco, California 94107. Phone: (415) 442–0220.

⋆ ⋆ ⋆  ⋆ ⋆ ⋆  ⋆ ⋆ ⋆

With ACTIV-L in place and linked to IGC’s Peacenet and Econet, the first stage of the PACH network we are still working to expand was attained, a network which, like Z and the rest of the alternative press, could help circumvent the media information blockade, as well as serve as a focal point for democratic participation, discussion, and action. This potential was tested during the critical months leading to and following the Gulf “War.”

ACTIV-L and IGC Confront the Gulf Crisis

As Z readers are well aware of, one of the (few) positive concomitants of the Gulf Crisis was a surge of activism and the possibility of building a national consciousness leading to a broad-based long-term movement or coalition.

Space for critical thinking and coalition-building also opened in the electronic domains (even as it was slammed shut for viewers of mainstream TV Nooz), and during the Gulf Crisis and subsequent Gulf War, there was a swelling in “online” activity and interest as well.

ACTIV-L in cooperation with IGC responded by distributing critical information and updates to activists, not readily available elsewhere, about what was happening and why. This included detailed reports on the destruction in Iraq, and analyses of the war and the U.S. media by commentators generally excluded from public view, including editorials from the Nation, In These Times (typed in by readers) and news releases and articles by FAIR (posted by FAIR to Peacenet and forwarded to ACTIV-L). ACTIV-L subscribers were also informed about protests and projects in Europe far more extensively than their TV-viewing counterparts.

There was also extensive discussion on ACTIV-L and interaction among subscribers, for example, soliciting specific information for a pamphlet or exchanging information among campus activists about what sorts of actions were taking place elsewhere, with what tactics, and with what success.

Besides the regular email subscribers to the list — by then some 1,000 people — many readers would download articles and print them out, distributing them to their local activist groups or using them for public education, in pamphlets, talks, and so on. The net effect ACTIV-L and IGC had during the Gulf Crisis we cannot measure, but it is my hope that a glimpse is provided in the accompanying side-bar by the sampling of electronic mail I received during that period.

########################################################################
#                            SideBar:                                  #
########################################################################
#                                                                      #
#               ==========================================             #
#               Electronic Activism during the Gulf Crisis             #
#               ==========================================             #
#                                                                      #
# >>SPREADING INFORMATION:<<                                           #
#                                                                      #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -     #
#                                                                      #
#     From: Michael Perelman <michael@ecst.csuchico.edu>               #
#     Subject: thanks                                                  #
#     To: harelb@cabot.dartmouth.edu                                   #
#     Date: Thu, 28 Feb 91 13:42:22 PST                                #
#     Mailer: Elm [revision: 64.9]                                     #
#                                                                      #
# The material that you post has been invaluable in preparing talks    #
# and spreading information about the war.  Thanks.  Michael Perelman  #
#                                                                      #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -     #
#                                                                      #
# >>MORAL SUPPORT AND SPURRING LOCAL GROUPS:<<                         #
#                                                                      #
#     Date: Sat, 19 Jan 91 18:29:05 -0500                              #
#     From: Carl Turner <turner@psych.ufl.edu>                         #
#     To: harelb@dartvax.dartmouth.edu                                 #
#     Subject: Thank you                                               #
#                                                                      #
# Thank you for the stories posted about the anti-war activities going #
# on around the country.  I pulled a bunch of them off alt.activism    #
# and read them at a meeting of the Middle East Peace Group here in    #
# Gainesville.  To the 60 or so people at the meeting, it was          #
# obviously a great feeling to know that they were part of a           #
# nationwide movement.  Really, people were clapping and cheering at   #
# some of the stories.                                                 #
#                                                                      #
# There seems to be a virtual news blackout in the media about         #
# anti-war activities, and the things that do get reported are grossly #
# misleading (e.g., the San Francisco demonstrations).  [...]          #
#                                                                      #
#         --Carl Turner, turner@webb.psych.ufl.edu                     #
#                                                                      #
########################################################################

########################################################################
#                                                                      #
#     Date: Wed, 30 Jan 91 17:04:15 pst                                #
#     From: marie weston <mariew@hprpcd.rose.hp.com>                   #
#     Subject: 2nd article on Gulf...                                  #
#                                                                      #
# Thank you for posting this speech by Phil Agee.  I found it very     #
# interesting and I even printed off copies for several people I know, #
# including my SO, who is an Air Force Pilot.  I don't think he        #
# appreciated it as much as I did but it gave him something to think   #
# about.                                                               #
#                                                                      #
# Thanks again,                                                        #
# Marie                                                                #
#                                                                      #
# ["SO" stands for "Significant Other", another common UseNet phrase   #
#  --HB]                                                               #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -     #
#                                                                      #
# >>SUPPORTING COMMUNITY EFFORTS:<<                                    #
#                                                                      #
#     [From morrisr@ucs.indiana.edu, Jan. 15, 1991]                    #
#     [Subject: praise for ACTIV-L                ]                    #
#                                                                      #
# Greetings from Indiana University in Bloomington.  Our activist      #
# community is growing steadily. Activ-L has contributed to this       #
# growth by helping to inform many folks of events concerning the      #
# Gulf, Central America, and other places around the US and world.  We #
# now have over 70 people getting some to the things that come in on   #
# ACTIV-L! Interest is high.                                           #
#                                                                      #
# I speak for all of us here that we really appreciate all the work    #
# you are putting in to this. Believe me, your efforts have helped     #
# us a lot. We would not have access to this info if you were not      #
# doing what you do.                                                   #
#                                                                      #
# In solidarity,                                                       #
#                                                                      #
# Roger Morris                                                         #
# Bloomington, IN                                                      #
#                                                                      #
########################################################################

########################################################################
#                                                                      #
# >>IN THE U.S.<< ...                                                  #
#                                                                      #
# Date: 5 Jan 91 16:16:00 CDT                                          #
# From: "Mary Lynne Dittmar" <uahmld01@asnuah.asn.net>                 #
# Subject: P.A.C.H.                                                    #
#                                                                      #
#   Harel:                                                             #
#                                                                      #
#    Hi.  My name is Mary Lynne Dittmar, I'm a faculty member at       #
# University of Alabama, Huntsville.  There is a busy, thriving        #
# activist community here, probably in response to the fact that the   #
# Redstone Arsenal and NASA (Marshall Space Flight Center) are also    #
# here.  ACTIV-L has been tremendously helpful; it has spurred on a    #
# local group of academics (we're currently putting together a series  #
# of public forum/lecture/discussion meetings on the Perisan Gulf),    #
# provided networking with other activists across the country, and     #
# generally provided emotional and spiritual support -- we are not,    #
# after all, alone; there is a focus and direction to what we do.  I   #
# just wanted to let you know this -- especially after having read     #
# your posting of 4 January about effectiveness.  Also, after reading  #
# it, I wanted to ask for the P.A.C.H. document which you indicated    #
# was available by e-mail [...]                                        #
#                                                                      #
#     BITNET:  RADICAL@UAHVAX1 or UAHMLD01@UAHVAX1                     #
#   Thanks,                                                            #
#      Mary Lynne                                                      #
#                                                                      #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -     #
#                                                                      #
# ... <<AND IN EUROPE:<<                                               #
#                                                                      #
#     Date:     Tue, 5 Mar 91 14:47:48 MET                             #
#     From: Stefan Merten <merten@uklirb.informatik.uni-kl.de>         #
#     To: harelb@arthur                                                #
#     Subject:  Your forwards from PeaceNet                            #
#                                                                      #
# Hi!                                                                  #
#                                                                      #
# First of all I want to thank you for forwarding all these very       #
# interesting (and little biased :-) ) articles to UseNet.  I          #
# extracted many of them.  Though the following holds true for the     #
# rest of the net, the best for me European is to get informations     #
# from inside USA this way.  A quite difficult thing for a normal      #
# person.                                                              #
#                                                                      #
# Don't stop to forward backgrounds about the war, I need them! For    #
# example parts of these informations are used in a series of          #
# flyers which present backgrounds of the conflict and which is        #
# continued after the war.                                             #
#                                                                      #
#                                                     Stefan           #
#                                                                      #
########################################################################

A New Usenet Newsgroup

This flurry of activity prompted us to create a new Usenet newsgroup, misc.activism.progressive (MAP), hoping to expand the project by reaching a wider audience.

This necessitated having an online “vote” — newsgroup-creation is the only facet of Usenet which follows a pattern of democracy rather than anarchy — and hence building up enough support to attain the necessary two-to-one margin over the significant minority of vocal (and paranoid) reactionaries opposed to anything “leftist”, let alone a “moderated” forum which we were requesting to host, with the specter of electronic indoctrination camps used often — an absurdity in light of the ability of any Usenet reader to press a key to read any of countless other political, un-moderated, and mainstream- or right-wing-dominated forums.

The lengthy process was an emotional strain which took its toll, yet today that mentality can be a source of amusement, paralleling paranoid Senators worrying about a “leftist” PBS and threatening to freeze funding — this in light of the documentation, recently circulated again by FAIR on Peacenet (and through our newsgroup and other Usenet newsgroups, to tens of thousands of online readers), showing the pervasive right-wing bias on “public” TV.

The newsgroup “passed”, but not before several more attempts where made to wreck the enterprise, including charges — actually, totally groundless assertions — about every conceivable type of impropriety or “fraud,” including unprecedented post-vote “surveys” to ask the voters listed as “in favor” whether they in fact voted this way. The voters replied in the affirmative, but not before voicing their disgust at the charges, including charges that the IGC directorship “faked” the crucial hundred-odd Peacenet “yes” votes. Judging by this quite hostile reaction, one might have concluded there and then that we were on to something good... (-:

MAP’s Mission

Aside from wanting to create a forum with a potentially larger audience, MAP was to become Usenet’s first genuinely activist group, since the existing alt.activism had been all but taken over by “conservatives” — reactionaries and general weenies bent on endless arguments, infantile sarcastic “jokes” and otherwise “flaming” any post which might lead to something constructive in the vague direction of that newsgroup’s name — “Boy, Saddam would really love your positions!” was (and is) par for the course.^{15}

This was possible because alt.activism is an “un-moderated” newsgroup, to which anyone could post a message (and usually did). MAP’s charter was for a moderated newsgroup which would:

provide news, information, and resources of interest to activists and anyone interested in promoting peace; economic, legal, and social justice and equality; a clean environment and safe workplace; democratic control over domestic and foreign policy, and other issues of concern to the progressive community.

and which “is not for political discussion, although we welcome inquiries and requests for information and references, announcements of local activist organizing...” — ACTIV-L would serve as a forum for discussion^{16} Together with forwards from the IGC networks, other mailing lists, and readers’ contributions, these resources include books (reading lists), listings of alternative publications, videos, media and government telephone numbers, announcements of speaking tours, and other data activists rely on.

15In fact, I recently received email from an old acquaintance which opened: “After a year without net.access, I’m pleased to see you’re still there. Looks like I missed the virtual abandonment of alt.activism to the sabateurs and the creation of misc.activism.progressive” 16ACTIV-L has since spawned ACTIV-D for discussions, while ACTIV-L now runs just the articles posted to MAP, for people who do not have access the Usenet.

#############################################################################
# [SIDE-BAR: "Spreading The Word"]                                          #
#############################################################################
#                  ===================================                      #
#                  How the Nets help ``Spread the Word''                    #
#                  ===================================                      #
#                                                                           #
# >>INFORMATION:<<                                                          #
#                                                                           #
# Date: Fri, 1 Feb 91 15:12:13 -0500                                        #
# From: painter@world.std.com (Alan J Painter)                              #
# Subject: FAIR report -- Re: The media and the war: unbiased coverage      #
#                 or rooting for the home team?                             #
#                                                                           #
# I simply wanted to thank you for the great information that you have      #
# been posting to alt.activism.  I have just started reading this           #
# group (in fact, just started reading news in general due to my            #
# previous unconnectedness) and I find that the raw information about       #
# El Salvador, statistical reportage and such are extremely useful.         #
# [...]                                                                     #
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
#         [Referring to an article we posted to the UseNet                  #
#         which came from a post to PeaceNet by Project Censored:]          #
#                                                                           #
# Date: Mon, 25 Feb 91 21:14:24 PST                                         #
# From: robert Arrignton <arringto@nevada.edu>                              #
# To: harelb@cabot.dartmouth.edu                                            #
# Subject: Re: TOP TEN CENSORED STORIES FOR 1990                            #
# Newsgroups: alt.activism,talk.politics.misc,                              #
#         alt.desert-storm,alt.censorship                                   #
# Organization: University of Nevada, System Computing Services             #
#                                                                           #
# VERY GOOD, INFORMATIVE POST.  THIS HAS GOT TO BE THE MOST                 #
# INFORMATIVE, UNDERRATED ARTICLE I'VE SEEN THIS MONTH.                     #
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
# Date: Sat, 1 Jun 91 05:33:25 MDT                                          #
# From: Harry Brooks <brooksh@jacobs.CS.ORST.EDU>                           #
# To: harel@dartvax.dartmouth.edu                                           #
# Subject: compliment                                                       #
# Cc: rfrancis@u.washington.edu                                             #
#                                                                           #
# Hi Harel,                                                                 #
# I just wanted to compliment you on howyou have listed                     #
# and grouped items in the "activ-l index activ-l" file:                    #
# it is very easy to read and to locate desired information.                #
# Also it communicates plainly.                                             #
# Keep up the good work!                                                    #
# harry                                                                     #
#                                                                           #
#############################################################################

#############################################################################
#                                                                           #
# Resent-From: "Rich Winkel UMC Math Department" <MATHRICH%UMCVMB.bitnet>   #
# Date:         Mon, 21 Jan 91 20:19:35 EST                                 #
# From: Susan Bramhall <SUSAN%YALEVM.bitnet>                                #
# Subject:      Re: Re: Are we still alive?                                 #
#                                                                           #
# Thought you'd want to see yet another bit of encouragement ...            #
# Rich                                                                      #
#                                                                           #
# P.S.  There seems to be a hangup on the net today, which has blocked      #
# most list subscribers from getting anything.  I've tracked it down        #
# and complained to the appropriate people.  Hopefully it will clear        #
# soon.                                                                     #
#  -------------------------- Original note follows-------------------------#-
# It turns out that the list was held at this end because the spool         #
# was over 40% full.  So I now have all the backed up mail.  I am           #
# sorry to have bothered you but I guess this tells you how incredibly      #
# valuable the list has become to me.  Many, many thanks for the time       #
# you put into it.                                                          #
#          Susan                                                            #
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
# >>ANALYSIS:<<                                                             #
#                                                                           #
# From: eiverson@NMSU.Edu                                                   #
# Date: Wed, 30 Jan 91 21:22:27 MST                                         #
# Subject: GULF NEWSpeak -- "Pundit Watch" by Greenpeace                    #
#                                                                           #
# Please forward my wholehearted praise to the authors of the above         #
# piece.  Excellent writing!                                                #
# -------------------------------------------------------------------       #
# Eric Iverson                            Internet: eiverson@nmsu.edu       #
# Computing Research Lab                                                    #
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
# >>VIDEOS:<<                                                               #
#                                                                           #
# Re: COVERUP video                                                         #
# Hi--                                                                      #
#                                                                           #
# I saw your recommendation of the video COVERUP: Behind the                #
# Iran-Contra Affair a few months ago and I finally got around to           #
# ordering it.  I had heard some of the stuff before, but it really         #
# fueled my discontent with the Bush Shadow Government.  Just wanted        #
# to thank you for recommending this video, and for your excellent          #
# posts on Usenet [...]                                                     #
#                                                                           #
# Freedom,                                                                  #
# Drew                                                                      #
#############################################################################

#############################################################################
#                                                                           #
#+---------------------+--------------------------------+-------------------#
#| Drew Davidson       | Disclaimer:  These are my own  |  JUST SAY KNOW!   #
#| davidson@cs.unc.edu | opinions, not those of UNC-CH. | HEMP FOR VICTORY! #
#+---------------------+--------------------------------+-------------------#
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
# >>The Alternative Press:<<                                                #
#                                                                           #
# Date: Thu, 20 Dec 90 16:58:02 -0500                                       #
# From: doog@travis.ssd.csd.harris.com (Doug Scofield)                      #
# To: harelb%arthur@gargoyle.uchicago.edu                                   #
# Subject: In These Times                                                   #
#                                                                           #
# harelb@arthur.uchicago.edu (Harel Barzilai) writes:                       #
# ) ++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++      #
# )                                                                         #
# ) For a free issue of the national weekly newspaper In These Times,       #
# ) just cut and mail their card, which is reproduced below... The          #
# ) address to put on the back of this "postcard" is:                       #
# )                                         In These Times                  #
# )                                         1912 Debs Avenue                #
# )                                         Mt. Morris, IL  61054           #
#                                                                           #
# I tried the sample copy of "In These Times" because of your               #
# signature, and i'm glad i did.  Finally i've found a newspaper            #
# that's got the slant i like, reporting on issues i care about,            #
# and a size that allows me to read it in an hour.  It's like               #
# a concentrated news pill.  Thanks!!                                       #
#                                                                           #
# doug scofield                    doog@ssd.csd.harris.com  (preferred)     #
# harris computer systems                  uunet!hcx1!doog                  #
#             "Nothin' left to do but :-) ;-) 8-)"                          #
#                                                                           #
#  - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - -          #
#                                                                           #
# Hi                                                                        #
#                                                                           #
# I am on the faculty at USC and I recently read your posting of Phil       #
# Agee's censored speech.  In the article you mentioned Z magazine.         #
# Could you tell me a little more about this magazine and how one can       #
# get a subscription.                                                       #
# Thank you.                                                                #
#                                                                           #
# Sarma Sastry                                                              #
# Dept. of EE-Systems                                                       #
# Univ. of So. Cal.                                                         #
# LA,CA 90089-0781                                                          #
# email : sastry@vishnu.usc.edu                                             #
#                                                                           #
#############################################################################

#############################################################################
#                                                                           #
# From: pape@agcrr.bio.ns.ca (Andrew Pape)                                  #
# Subject: PEACENET                                                         #
# To: harelb@cabot.dartmouth.edu                                            #
# Date: Sun, 19 May 91 21:48:14 ADT                                         #
# X-Mailer: ELM [version 2.3 PL11]                                          #
#                                                                           #
# Hi                                                                        #
#                                                                           #
# I wrote you a letter last week to find out more about Peacenet.  It       #
# looks incredible!  I want to join.  Thank you for sending me              #
# information on it.                                                        #
#                                                                           #
#############################################################################

(Part (II) continues with “MAP^{17} Today and the PACH Project”)

Harel Barzilai is a graduate student at Cornell University. He is co-founder and co-moderator of the ACTIV-L mailing list and UseNet’s misc.activism.progressive (MAP), and an avid electronic activist, currently working on the PSN-phase (Electronically linked Progressive Student Newspapers) of the PACH project, described in Part II of this article.

17Currently with over 12,000 readers

* Typesetting by L^{A}_{M}S-TEX